the second side of a triangle is 2dm less than twice the first side and the third side is 4dm less than twice the second side. how long is each side if the perimeter of the triangle is 18dm.
Answer by mananth(16949)

let first side be x dm
second side = 2x-2
third side =2(2x-2)-4
..
x+2x-2+4x-4-4 =18
7x-10=18
7x = 28
x= 4 first side
2x-2 = 8-2 = 6 second side
third side = 2(2*4-2)-4=8
..
4+6+8 =18dm